Two equally charged insulating balls each weigh 0.10 g and hang from a common point by identical threads 30 cm long. The balls repel each other so that the separation between their centers is 6.2 cm. What is the magnitude of the charge on each ball?
Answer #1

θιθ Tcose L.-30cm Tsine Ct d-6.2cm mg

First find heta

sinθ = (d/2) / (30cm)

sin (3.1cm)/(30cm)

1 θ sin- (3.1/30)

heta =5.93^{circ}

=============

Consider all vertical Forces

sum F_{y}=0

cOS

mg / cose

T = (1 * 10-4kg)(9.81)/cos593

T- 9.86 * 10-4V

=============

Consider all Horizontal Forces

Fr = 0

F=Tsin heta

kq^{2}/d^{2}=Tsin5.93

q^{2}=d^{2}*Tsin5.93/k

q^{2}=(0.062m)^{2}*(9.86*10^{-4}N)sin5.93/(9*10^{9})

q=6.596*10^{-9}C

ANSWER: q=6.596nC

=====================